Calculate the percent yield using data belowQuestion 1. Following the Task 1 experimental procedure as per the Project 10 Guide (pages 5 – 7), 1.993 g of coconut oil is
Answer #1

Reaction involved

TAG + 3MOH ------> 3(M salt of fatty acid) + Glycerol

Here, M is Sodium (Na) or Potassium​​​​​​ (K)

Moles of coconut oil taken = 1.993 g/(869.4 g/mol)

= 0.002292 moles

Theoretical moles of soap to be formed = 3*0.002292 = 0.006876 moles

Theoretical mass of sodium soap to be formed from coconut oil = 0.006876*299.1 = 2.057 g

Mass of soap formed = 0.817 g

% yield = (0.817/2.057) x 100 = 39.72 %

In case potassium soap is formed:

Theoretical mass of potassium soap from coconut oil = 0.006876*315.2 = 2.167 g

% yield = (0.817/2.167)*100 = 37.70 %